OpenStudy (ka.wehi):

A triangular piece of rubber is stretched equally on all three sides (without distorting or changing its shape) such that the new sides are 2 times the length of the original sides. How does the area of the triangle change?

jimthompson5910 (jim_thompson5910):

Each side is doubled, so the area is multiplied by 4. Why 4? because 2^2 = 4

jimthompson5910 (jim_thompson5910):

If each side is tripled, then the area would increase by a factor of 9 3^2 = 9

jimthompson5910 (jim_thompson5910):

If each side is quadrupled, then the area would increase by a factor of 16 4^2 = 16

jimthompson5910 (jim_thompson5910):

and so on...

OpenStudy (ka.wehi):

If ADE ABC in the diagram, which pair of ratios must necessarily be equal?